OpenStudy (anonymous):

For the following quadratic function, name the "a" value, the "h" value, and the "k" value. y=2(x-8)^2+10 A. a=2, h=8, k=10 B. a=2, h= -8, k = -10 C. a = 8, h = 10, k = 2 D. a = 10, h = 2, k = 8

jigglypuff314 (jigglypuff314):

referring to the form of y = a(x - h)^2 + k so for y = 2(x - 8)^2+10

jigglypuff314 (jigglypuff314):

a = 2, h = -8, k = 10